What do you understand by the term “closed economy”? For a closed economy show, 

National savings = National Investment


1
Expert's answer
2021-12-09T09:27:26-0500

The closed economies are those economies who do not get indulged in international trade with the nations around the globe. 

These closed economies neither buy foreign goods, and services, nor sell their domestically produced goods and services to the nations. 

 

In the closed economies, the amount of investments are equal to the amount of savings. 

 "income=consumption+investiment+government \space spending\\Y=C+I+G\\I=Y-C-G\\national \space saving=private\space saving+public\space saving\\S=Y-taxes-consumption+(taxes -government \space spending)\\S=Y-T-C+T-G\\S=Y-C-G\\S=I"


Hence, we can say that savings are equal to investment in closed economy.